# AuxDet: Auxiliary Metadata Matters for Omni-Domain Infrared Small Target Detection

### v1

@@ Abstract
Omni-domain infrared small target detection (IRSTD) poses formidable challenges, as a single model must seamlessly adapt to diverse imaging systems, varying resolutions, and multiple spectral bands simultaneously. Current approaches predominantly rely on visual-only modeling paradigms that not only struggle with complex background interference and inherently scarce target features, but also exhibit limited generalization capabilities across complex omni-scene environments where significant domain shifts and appearance variations occur. In this work, we reveal a critical oversight in existing paradigms: the neglect of readily available auxiliary metadata describing imaging parameters and acquisition conditions, such as spectral bands, sensor platforms, resolution, and observation perspectives. To address this limitation, we propose the Auxiliary Metadata Driven Infrared Small Target Detector (AuxDet), a novel multi-modal framework that fundamentally reimagines the IRSTD paradigm by incorporating textual metadata for scene-aware optimization. Through a high-dimensional fusion module based on multi-layer perceptrons (MLPs), AuxDet dynamically integrates metadata semantics with visual features, guiding adaptive representation learning for each individual sample. Additionally, we design a lightweight prior-initialized enhancement module using 1D convolutional blocks to further refine fused features and recover fine-grained target cues. Extensive experiments on the challenging WideIRSTD-Full benchmark demonstrate that AuxDet consistently outperforms state-of-the-art methods, validating the critical role of auxiliary information in improving robustness and accuracy in omni-domain IRSTD tasks.
1 Introduction
@@
Infrared imaging offers unparalleled advantages in capturing target information under adverse weather conditions and limited visibility scenarios, which has enabled wide-ranging applications in marine resource utilization (Han et al., 2022), precision navigation (Sun et al., 2021b), and ecological monitoring (Hao et al., 2024). However, due to the long-range imaging distance, targets of interest typically exhibit extremely limited spatial resolution, lack texture and structure, and often blend into complex backgrounds (Zhao et al., 2022), making them easily overwhelmed by environmental clutter. Therefore, infrared small target detection (IRSTD) still remains a challenging task for computer vision.
@@
Existing IRSTD methods primarily focus on intra-image feature extraction to distinguish dim targets from complex backgrounds (Chen et al., 2013). Recently, with the rise of deep learning, many efforts have attempted to address IRSTD through neural architectures that emphasize robust feature extraction and refined attention mechanisms (Chi et al., 2025). Representative works like ALCNet (Dai et al., 2021b) introduced attention-guided local contrast mechanisms to amplify subtle target-background discrepancies, while ISNet (Zhang et al., 2022b) developed infrared-specific spatial transformers for noise-robust feature learning. Despite these advances, existing methods exhibit two critical shortcomings when scaling to real-world scenarios:
  1. 1.
@@
    Over-Specialization: Most approaches optimize for narrow operational domains, such as land-based platforms with LWIR bands or space-based SWIR imagery (Li et al., 2024), failing to generalize across the sensor-platform-spectrum continuum inherent to omni-domain applications.
    Figure 1: Multidimensional Metadata Analysis of WideIRSTD-Full Dataset [^1] https://github.com/XinyiYing/WideIRSTD-Dataset. (a) Target-background contrast divergence reflecting platform-specific target heterogeneity; (b) Platform-band joint distribution characteristics; (c) Cross-platform resolution distribution (log-scale) revealing multi-scale detection requirements; Key findings: 1) Multi-source metadata can provide physically interpretable guidance for detection tasks; 2) Single-category metadata exhibits representational limitations, while multi-category joint modeling can offer comprehensive auxiliary support for image processing/detection.
  2. 2.
@@
    Over-Engineering: Researchers increasingly rely on domain-specific techniques to enhance detection accuracy, such as pre-identifying sea regions through background decomposition for maritime targets (Hu et al., 2025). While these methods reduce task complexity in narrow domains, they inherently compromise cross-domain generalization capabilities.
@@
To meet the growing demand for large-scale, multi-platform, and real-world deployable detection systems, in this paper, we formalize a new task, Omni-Domain Infrared Small Target Detection (Omni-IRSTD). As shown in Fig. 1, this task aims to unify the IRSTD paradigm across heterogeneous sensing domains by addressing four core dimensions: (1) multi-platform observations (e.g., land-based, aerial-based, and space-based), (2) multi-target types (point-like, spot-like, and extended targets), (3) multi-spectral bands (e.g., LWIR, NIR, SWIR), and (4) multi-resolution imaging (from $256\times 256$ to $6000\times 6000$(Li et al., 2024). Compared to traditional specific-domain IRSTD, Omni-IRSTD presents substantial challenges due to dramatic variations in environmental contexts, background characteristics, and target appearance distributions. These cross-domain disparities significantly impact model generalization capabilities, as evidenced in Tab. 1, where state-of-the-art methods (Sun et al., 2023a; Yang et al., 2024) experience catastrophic performance degradation on omni-domain benchmarks. This performance gap underscores the need for a paradigm shift to address the challenges posed by the diverse and complex nature of Omni-IRSTD.
@@
We argue that the fundamental limitation of existing approaches stems from their paradoxical pursuit of unified feature representation across physically divergent imaging domains. When naively aggregating multi-domain training data, the gradient directions from heterogeneous domains exhibit severe conflicts in parameter space forcing the network into suboptimal equilibrium states that compromise all domains (Zhao et al., 2022). This observation sparks a natural insight: since the ideal feature representations differ across domains is neither practical nor effective, instead, the model should be aware of which domain a given input belongs to, and adapt its representation strategy accordingly. However, conventional single-modality visual models lack the capacity to infer domain-specific imaging conditions solely from raw pixel data (Zhang et al., 2025). This raises a critical question: can a model know the domain of an input sample in advance to enable such adaptability?
@@
Our answer is yes. In practical infrared sensing systems, each captured image is inherently accompanied by a set of auxiliary metadata that records the physical conditions under which the image was acquired, such as sensor platform (e.g., airborne, satellite), spectral band (e.g., LWIR, NWIR), spatial resolution, and observation angle. Rather than being incidental or redundant, these metadata descriptors serve as compact, semantic-rich representations of the domain to which each sample belongs (Zhu et al., 2024). In essence, metadata offers a low-dimensional yet highly informative abstraction of the imaging domain, encoding the very factors that drive domain-specific feature distributions and appearance patterns. By explicitly incorporating this metadata into the learning process, we can transform the detection paradigm from a passive, vision-only modeling regime to an active, context-aware framework, where the model is empowered to adapt its representation strategy conditioned on the domain characteristics of each input sample (Gao et al., 2025).
@@
Building upon these insights, we propose the Auxiliary Metadata Driven Detector (AuxDet), a novel framework for Omni-IRSTD that redefines the conventional visual-only paradigm by leveraging auxiliary metadata for multi-modal scene-aware optimization. Unlike existing methods, AuxDet dynamically integrates metadata, such as spectral bands, sensor parameters, and observation conditions, with visual features through a carefully designed Multi-Modal Dynamic Modulation (M2DM) module. This module employs a multi-layer perceptron (MLP)-based fusion strategy to embed metadata semantics into the feature space, enabling sample-specific modulation that enhances target discriminability amid complex backgrounds. By incorporating cross-layer differential features, M2DM further compensates for the scarcity of small target cues, fostering robust contextual representations. Further, to address the trade-off between high-level semantics and fine-grained target details, we propose a Lightweight Edge Enhancement Module (LEEM), which utilizes 1D convolutional blocks initialized with local differential priors to refine edge responses and boost detection precision.
Extensive experiments on the challenging WideIRSTD-Full benchmark (Li et al., 2024) demonstrate the effectiveness of our approach. Notably, compared to visual-only baselines, AuxDet yields over 30% improvement in detection accuracy, highlighting the critical role of auxiliary metadata in enabling robust and generalizable omni-domain IRSTD.
The main contributions of this work are summarized as follows:
@@
  • We are the first to introduce auxiliary metadata into the IRSTD task, transforming the traditional vision-only detection paradigm into a multi-modal, scene-aware framework.
  • We propose the M2DM module that dynamically modulates visual features using metadata.
@@
  • We present the LEEM module, an ultra-lightweight module that amplifies edge responses essential for detecting extremely small targets.
2 RELATED WORK
2.1 Infrared Small Target Detection
IRSTD aims to detect objects that occupy only a few pixels against complex backgrounds, often under extremely low signal-to-noise ratio conditions. Early IRSTD methods primarily relied on handcrafted features and heuristic strategies. Filtering-based techniques (Yang et al., 2020; Xi et al., 2025) aimed to enhance target signals by suppressing background noise. Similarly, background modeling approaches, including temporal profile modeling (Du et al., 2022; Zhou et al., 2023) and low-rank and sparse decomposition (Gao et al., 2013; Dai et al., 2016; Liu et al., 2022b), sought to separate targets from their surroundings. While effective in controlled settings, these approaches lack generalization across diverse imaging scenarios where background variability and noise overwhelm their simplistic priors.
In recent years, deep learning methods have demonstrated remarkable success in learning robust feature representations directly from infrared images. Convolutional neural networks (CNNs), such as ALCNet (Dai et al., 2021b) and ISNet (Zhang et al., 2022b), harness multi-scale feature extraction and attention mechanisms to amplify target saliency while suppressing background interference. More recently, Transformer-based architectures (Wang et al., 2022; Pan et al., 2023; Yuan et al., 2024) have leveraged long-range dependencies to further enhance detection accuracy. SCTransNet (Yuan et al., 2024) leverages spatial-channel cross transformer blocks on top of long-range skip connections to address the challenge of high similarities between target and background. However, as revealed in our analysis, these models are trained on datasets of specific application scenarios (Zhang et al., 2022b; Dai et al., 2021a, 2023), which suffer catastrophic performance drops when deployed to unseen domains due to domain-specific feature overfitting. Moreover, omni-domain IRSTD task amplifies this critical limitation, demanding a single model to perform reliably across multiple imaging systems and conditions.
@@ 2.1 Infrared Small Target Detection
In contrast to previous works, our work introduces a fundamental shift in both task formulation and modeling paradigm. We position AuxDet as the first framework to explicitly tackle omni-domain IRSTD through metadata-guided multi-modal learning:
@@
2.2 Metadata-Enhanced Visual Recognition
@@
Auxiliary metadata, such as geographic location, observation time, and viewing angle, has proven valuable in enhancing semantic understanding in image classification tasks. These contextual cues provide high-level priors about environmental conditions, lighting, and scene composition, which are often difficult to infer from raw pixel data alone. Tang et al. (Tang et al., 2015) first demonstrated the value of GPS data in improving classification accuracy, leveraging location-specific priors to disambiguate visually similar categories. Chu et al. (Chu et al., 2019) integrated temporal metadata and viewing angles via multi-layer perceptrons (MLPs) to disentangle seasonal variations and perspective distortions. However, these early methods (Mac Aodha et al., 2019; Terry et al., 2020) relied on simplistic fusion strategies, limiting their ability to capture complex interactions between visual and auxiliary data.
@@
To overcome these shortcomings, Yang et al. (Yang et al., 2022) proposed a dynamic MLP approach, using instance-specific weighted projections to map similar image features into distinct feature space regions, thus boosting classification precision. Xu et al. (Xu et al., 2024) advanced this further with a multi-stage fusion strategy, integrating early-stage input-level fusion with late-stage dynamic MLP structures. Gao et al. (Gao et al., 2025) emphasized early-stage fusion to retain fine-grained geographic details, striking a balance between computational efficiency and contextual richness. These developments highlight auxiliary information as a source of high-level semantic priors.
@@
However, these approaches are confined to image-level classification, and have not been extended to detection tasks, especially not to infrared small target detection. To the best of our knowledge, our work is the first to introduce and systematically model metadata as an auxiliary signal in the IRSTD domain.
@@
Figure 2: Overall architecture of AuxDet. Infrared images are processed through a backbone network to extract hierarchical features, while auxiliary data undergo preprocessing and fusion via MLPs. The two modalities enter into M2DM to achieve dynamic feature interaction and instance-level spatial-channel modulation, obtaining modulated features that are subsequently refined by LEEM. The optimized features are fused through the FPN and fed into the detection head for target localization.
3 METHOD
3.1 Overall Architecture

### v2

@@ Abstract
Omni-domain infrared small target detection (Omni-IRSTD) poses formidable challenges, as a single model must seamlessly adapt to diverse imaging systems, varying resolutions, and multiple spectral bands simultaneously. Current approaches predominantly rely on visual-only modeling paradigms that not only struggle with complex background interference and inherently scarce target features, but also exhibit limited generalization capabilities across complex omni-scene environments where significant domain shifts and appearance variations occur. In this work, we reveal a critical oversight in existing paradigms: the neglect of readily available auxiliary metadata describing imaging parameters and acquisition conditions, such as spectral bands, sensor platforms, resolution, and observation perspectives. To address this limitation, we propose the Auxiliary Metadata Driven Infrared Small Target Detector (AuxDet), a novel multimodal framework that is the first to incorporate metadata into the IRSTD paradigm for scene-aware optimization. Through a high-dimensional fusion module based on multi-layer perceptrons (MLPs), AuxDet dynamically integrates metadata semantics with visual features, guiding adaptive representation learning for each individual sample. Additionally, we design a lightweight prior-initialized enhancement module using 1D convolutional blocks to further refine fused features and recover fine-grained target cues. Extensive experiments on the challenging WideIRSTD-Full benchmark demonstrate that AuxDet consistently outperforms state-of-the-art methods, validating the critical role of auxiliary information in improving robustness and accuracy in omni-domain IRSTD tasks. Code is available at https://github.com/GrokCV/AuxDet.
I Introduction
@@
Infrared imaging offers unparalleled advantages in capturing target information under adverse weather conditions and limited visibility scenarios, which has enabled wide-ranging applications in marine resource utilization [1], precision navigation [2], and ecological monitoring [3]. However, due to the long-range imaging distance, targets of interest typically exhibit extremely limited spatial resolution, lack texture and structure, and often blend into complex backgrounds [4], making them easily overwhelmed by environmental clutter. Therefore, infrared small target detection (IRSTD) still remains a challenging task for computer vision.
Figure 1: Traditional vs. unified detection. AuxDet leverages metadata for effective omni-domain generalization using one unified model.
One of the major difficulties in infrared small target detection lies in the extremely small target size and the lack of distinguishable textures or colors, making them prone to being submerged in background noise. Traditional detection methods rely on prior assumptions about targets and backgrounds, using mathematical models for extraction [5, 6, 7, 8, 9, 10, 11, 12, 13]. However, as background complexity increases, these methods struggle with robustness and adaptability.
@@
Recently, with the rise of deep learning, many efforts have attempted to address IRSTD through neural architectures that emphasize robust feature extraction and refined attention mechanisms [14]. Representative works like ALCNet [15] introduced attention-guided local contrast mechanisms to amplify subtle target-background discrepancies, while ISNet [16] developed infrared-specific spatial transformers for noise-robust feature learning. Despite these advances, existing methods exhibit two critical shortcomings when scaling to real-world scenarios:
@@
  • Over-Specialization: Most approaches optimize for narrow domains, such as land-based Long-Wave Infrared (LWIR) or space-based Short-Wave Infrared (SWIR) imagery, failing to generalize across the sensor-platform-spectrum continuum of omni-domain applications.
@@
  • Over-Engineering: Researchers increasingly rely on domain-specific techniques to enhance detection accuracy, such as pre-identifying sea regions through background decomposition for maritime targets [17]. While these methods reduce task complexity in narrow domains, they inherently compromise cross-domain generalization capabilities.
Figure 2: The benefit of leveraging auxiliary information to assist infrared small target detection lies in its ability to guide the network towards inference objectives that typically require deep architectures, but at an extremely low computational cost.
Figure 3: Multidimensional distribution characteristics of the WideIRSTD-Full Dataset. (a) Cross-platform target-background contrast: Resolution-normalized visualization of platform-specific heterogeneity in background complexity and small target morphology. (b) Cross-platform resolution distribution: x/y-axes represent log-transformed image height and width ranges, revealing resolution disparities across observation platforms (Air/Land/Space). (c) Platform-band correlation chord diagram: Coupling relationships between observation platforms (Air/Land/Space) and infrared bands (Long-Wave Infrared, LWIR; Near-Infrared, NIR; Short-Wave Infrared, SWIR). Key findings include: (1) Inter-platform resolution intersection: Significant overlapping regions between Air (green) and Land (blue) platforms necessitate cross-platform multi-scale detection; (2) Platform-dependent target characteristics: Distinct platform-specific distributions of target resolutions and categories induce varying detection difficulties; (3) Spectral-platform association rules: Space-based observations (orange) are NIR-dominant, while aerial/terrestrial platforms (green/blue) solely utilize LWIR.
@@
To meet the growing demand for large-scale, multi-platform, and real-world deployable detection systems, in this paper, we formalize a new task, Omni-Domain Infrared Small Target Detection (Omni-IRSTD). As shown in Fig. 1 and Fig. 3, this task aims to unify the IRSTD paradigm across heterogeneous sensing domains by addressing four core dimensions: (1) multi-platform observations (e.g., land-based, aerial-based, space-based), (2) multi-target types (point-like, spot-like, extended targets), (3) multi-spectral bands (e.g., LWIR, Near-Infrared (NIR), SWIR), and (4) multi-resolution imaging (from $256\times 256$ to $6000\times 6000$[18]. Compared to traditional specific-domain IRSTD, Omni-IRSTD presents substantial challenges due to dramatic variations in environmental contexts, background characteristics, and target appearance distributions. These cross-domain disparities significantly impact model generalization capabilities, as evidenced in Tab. I, where state-of-the-art (SOTA) methods experience catastrophic performance degradation on omni-domain benchmarks. This performance gap underscores the need for a paradigm shift to address the challenges posed by the diverse and complex nature of Omni-IRSTD.
@@
We argue that the fundamental limitation of existing approaches stems from their paradoxical pursuit of unified feature representation across physically divergent imaging domains. When naively aggregating multi-domain training data, the gradient directions from heterogeneous domains exhibit severe conflicts in parameter space forcing the network into suboptimal equilibrium states that compromise all domains [4]. This observation sparks a natural insight: since the ideal feature representations differ across domains is neither practical nor effective, instead, the model should be aware of which domain a given input belongs to, and adapt its representation strategy accordingly. However, conventional single-modality visual models lack the capacity to infer domain-specific imaging conditions solely from raw pixel data [19]. This raises a critical question: can a model know the domain of an input sample in advance to enable such adaptability?
@@
Our answer is yes. In practical infrared sensing systems, each captured image is inherently accompanied by a set of auxiliary metadata that records the physical conditions under which the image was acquired, such as sensor platform, spectral band, resolution, and observation angle. Rather than being incidental or redundant, these metadata descriptors serve as compact, semantic-rich representations of the domain to which each sample belongs [20]. In essence, metadata offers a low-dimensional yet highly informative abstraction of the imaging domain, encoding the very factors that drive domain-specific feature distributions and appearance patterns. By explicitly incorporating this metadata into the learning process, as shown in Fig. 1, we can transform the detection paradigm from a passive, vision-only modeling regime to an active, context-aware framework, where the model is empowered to adapt its representation strategy conditioned on the domain characteristics of each input sample [21].
More concretely, these metadata elements encode domain-specific priors through three key dimensions, as illustrated in Fig. 2:
  • Observation platform (e.g., air-based, space-based, or land-based) inherently conveys background complexity and target distribution characteristics through platform-specific attributes (e.g., in space-based remote sensing, the probability of certain geometric features in maritime targets increases significantly). At the same time, this information implicitly encodes atmospheric transmittance, aiding in the estimation of true radiative intensity. Consequently, during target detection, it guides the model to adopt more precise and adaptive processing strategies based on the observation source.
  • Spectral band of image acquisition(e.g., LWIR, SWIR, or NIR) serves as a distinctive signature of electromagnetic radiation. Variations in physical mechanisms governing different wavelength reflections directly determine the formation of target-background contrast and the statistical properties of noise distribution.
  • Image resolution acts as a crucial prior, enabling the model to infer the actual target size based on scaling factors, thus providing more refined decision support for detection tasks.
The combination of just a few characters representing these metadata elements offers highly valuable "differentiation descriptors" and "methodological guidance" for both images and networks, thereby circumventing the need for complex model design and structural optimizations in omni-domain infrared small target detection scenarios.
@@
Building upon these insights, we propose the Auxiliary Metadata Driven Detector (AuxDet), a novel framework for Omni-IRSTD that redefines the conventional visual-only paradigm by leveraging auxiliary metadata for multimodal scene-aware optimization. Unlike existing methods, AuxDet dynamically integrates metadata, such as spectral bands, sensor parameters, and observation conditions, with visual features through a carefully designed Multi-Modal Dynamic Modulation (M2DM) module. This module employs a multi-layer perceptron (MLP)-based fusion strategy to embed metadata semantics into the feature space, enabling sample-specific modulation that enhances target discriminability amid complex backgrounds. By incorporating cross-layer differential features, M2DM further compensates for the scarcity of small target cues, fostering robust contextual representations. Further, to address the trade-off between high-level semantics and fine-grained target details, we propose a Lightweight Edge Enhancement Module (LEEM), which utilizes 1D convolutional blocks initialized with local differential priors to refine edge responses and boost detection precision.
We summarize the main contributions of the paper as follows:
@@
  • We are the first to introduce metadata into the IRSTD task, transforming the traditional vision-only detection paradigm into a multimodal, scene-aware framework.
  • We propose the M2DM module, which injects auxiliary metadata into both channel-wise and spatial modulation processes. This design supports sample-specific adaptation of feature representations, empowering the detector with dynamic adaptability across complex and diverse environments.
@@
  • We design the LEEM, an ultra-lightweight module that amplifies edge responses essential for detecting extremely small targets.
Extensive experiments on the challenging WideIRSTD-Full benchmark [18] demonstrate the effectiveness of our approach. Notably, compared to visual-only baselines, AuxDet yields over 30% improvement in detection accuracy, highlighting the critical role of metadata in enabling robust and generalizable omni-domain IRSTD.
II Related Work
II-A Infrared Small Target Detection
Traditional infrared small target detection methods mainly rely on the characteristics of targets and backgrounds for target extraction. Depending on the prior knowledge or theoretical basis, these methods can be classified into: 1) methods based on local information, such as morphological Top-Hat transforms [5, 6] and local contrast metrics inspired by the human visual system [7, 8]. However, the characteristics of small targets make it difficult to match traditional structural elements, and the local contrast limits the utilization of global information, rendering these methods unable to accurately differentiate the subtle differences between targets and backgrounds. 2) Filtering-based methods, including spatial domain [9, 10] and frequency domain [11] filtering techniques, which fail to effectively separate targets from backgrounds when the spatial characteristics of the background and target are similar, and the forward and reverse transformations require considerable computational resources. 3) Low-rank representation-based methods, such as the Infrared Patch-Imag model [12] and Reweighted Infrared Patch-Tensor model [13], which assume that the background image exhibits low-rank characteristics when high-frequency components, such as random noise, are sparse. However, in practical applications, infrared image backgrounds are often more complex, and the low-rank assumption does not hold. While effective in controlled settings, these approaches lack generalization across diverse imaging scenarios where background variability and noise overwhelm their simplistic priors.
With the rapid development of deep learning and the release of various public datasets, deep learning-based detection methods have gradually emerged as the mainstream. These methods leverage the powerful modeling ability of deep neural networks to automatically learn features related to infrared small targets. Existing methods mainly focus on exploring network architectures, utilizing contrastive information, and emphasizing edge and shape information [22].
In terms of network architecture, Dai et al. [23] proposed an asymmetric context modulation mechanism to effectively solve the issue of scale mismatch between infrared small targets and objects in general datasets by employing complementary top-down and bottom-up modulation paths. Wu et al. [24] proposed an interactive cross-attention nested U-Net network to achieve multi-level target representation learning. Dai et al. [25] iteratively refined target bounding boxes through top-down cascaded detection heads. To establish long-range dependencies, Pan et al. [26] and Yuan et al. [27] introduced the visual Transformer architecture into infrared weak small target detection. Zhang et al. [28] introduced the Arbitrary Segmentation Everything model into the IRSTD domain, demonstrating outstanding performance.
With reference to contrastive information utilization, Dai et al. [15] integrated local contrast priors into convolutional networks to make full use of domain knowledge. Yu et al. [29] introduced a local contrast learning structure, inspired by non-deep learning methods, reducing the dependence on dataset sample size. Hu et al. [17] used a two-layer sliding window structure in their scene semantic extractor, using a larger window to compute background pixel changes and a smaller window to capture local contrast for the target, minimizing false alarms and assessing the likelihood of target appearance. Wang et al. [30] fused local contrast enhancement knowledge into the convolution operator design, effectively mitigating the blind nature and local extrema problem associated with autonomous learning in deep neural networks.
Regarding edge and shape information, Zhang et al. [16] proposed a new infrared shape network (ISNet), designed with edge blocks inspired by Taylor’s finite difference to extract useful edge features for predicting the precise shape of targets. Lin et al. [31] used central difference convolution and large kernel convolutions to extract discriminative representations of infrared small targets. Shi et al. [32] argued that the details and edges of tiny objects correspond to high-frequency components in the image spectrum and incorporated a high-frequency perception module in the lateral connections of the feature pyramid to generate high-frequency responses, enriching small target information in the backbone network features.
However, these models are trained on datasets of specific application scenarios [16, 23], which suffer catastrophic performance drops when deployed to unseen domains due to domain-specific feature overfitting. Moreover, omni-domain IRSTD task amplifies this critical limitation, demanding a single model to perform reliably across multiple imaging systems and conditions.
@@ II-A Infrared Small Target Detection
In contrast to previous works, our work introduces a fundamental shift in both task formulation and modeling paradigm. We position AuxDet as the first framework to explicitly tackle omni-domain IRSTD through metadata-guided multimodal learning:
@@
II-B Metadata-Enhanced Visual Recognition
@@
Auxiliary metadata, such as geographic location, observation time, and viewing angle, has proven valuable in enhancing semantic understanding in image classification tasks. These contextual cues provide high-level priors about environmental conditions, lighting, and scene composition, which are often difficult to infer from raw pixel data alone. Tang et al. [33] first demonstrated the value of GPS data in improving classification accuracy, leveraging location-specific priors to disambiguate visually similar categories. Chu et al. [34] integrated temporal metadata and viewing angles via multi-layer perceptrons (MLPs) to disentangle seasonal variations and perspective distortions. However, these early methods relied on simplistic fusion strategies, limiting their ability to capture complex interactions between visual and auxiliary data.
@@
To overcome these shortcomings, Yang et al. [35] proposed a dynamic MLP approach, using instance-specific weighted projections to map similar image features into distinct feature space regions, thus boosting classification precision. Xu et al. [36] advanced this further with a multi-stage fusion strategy, integrating early-stage input-level fusion with late-stage dynamic MLP structures. Gao et al. [21] emphasized early-stage fusion to retain fine-grained geographic details, striking a balance between computational efficiency and contextual richness. These developments highlight auxiliary information as a source of high-level semantic priors.
@@
It is clear that auxiliary information can explicitly introduce image-related features, which are often crucial high-level semantic information in many tasks. However, in the field of infrared small target detection, there has been no relevant application. To the best of our knowledge, our work is the first to introduce and systematically model metadata as an auxiliary signal in the IRSTD domain.
@@
Figure 4: Overall architecture of AuxDet. Infrared images are processed through a backbone network to extract hierarchical features, while auxiliary metadata undergo pre-encoding and fusion via MLPs. The two modalities enter into M2DM module to achieve dynamic feature interaction and instance-level spatial-channel modulation, obtaining modulated features that are subsequently refined by LEEM. The optimized features are fused through the Feature Pyramid Network (FPN) and fed into the detection head for target localization.
III Method
III-A Overall Architecture